# Distinguishing Neighborhood Representations Through Reverse Process of GNNs for Heterophilic Graphs

### v1

@@ 4 Experiments
The experimental section focuses on validating two research questions: 1) Can the reverse process produce sharpened representations? 2) Does the reverse process alleviate over-smoothing problems, enabling the construction of deeper layers?
Throughout this section, we denote models with additional reverse layers by ReP (Reverse Process). For example, GCN$+$ReP indicates the GCN backbone with the reverse process. We adopt weight sharing approach of GRAND-l for all experiments.
@@ 4.1.2 Experimental Setup and Baselines
We benchmark several neural architectures as baselines, including classic GNN models like GCN (Kipf & Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), and Graph Transformer (GT) (Shi et al., 2020) for more complex attention mechanisms. These baselines are augmented with skip connections and layer normalization. In addition, modifications proposed in Zhu et al. (2020) are made to GAT and GT, resulting in GAT-sep and GT-sep models. For heterophily-specific models, we use 10 models including H2GCN (Zhu et al., 2020), CPGNN (Zhu et al., 2021), GPR-GNN (Chien et al., 2020), FSGNN (Maurya et al., 2022), GloGNN (Li et al., 2022), FAGCN (Bo et al., 2021), GBK-GNN (Du et al., 2022), JacobiConv (Wang & Zhang, 2022), LRGNN (Liang et al., 2023), and OGNN (Song et al., 2023).
@@
For all experiments, we set the number of epochs to 1,000 and apply early stopping when there is no performance improvement for 100 consecutive epochs. For GRAND+ReP, we validate the hyperparameters that maximize the validation metric in the following ranges: learning rate $\in[10^{-5},10^{-1}]$, $T_{D},T_{R}\in[0,10]$, $d\in\{16,32,64,128,256,512\}$, $K\in[1,8]$, $d^{\prime}\in\{4,8,16,32,64,128\}$. With the Euler or Runge-Kutta methods, we search the step size over $[0.5,8]$ and tolerance scale over $[1,20000]$ with the Dormand-Prince method. For GCN+ReP and GAT+ReP, we validate the hyperparameters in the following ranges: learning rate$\in[10^{-5},10^{-1}]$, numbers of forward and reverse layers $L_{D},L_{R}\in\{1,2,4,8,16,32,64,128,256,512\}$, dropout probability $\in[0,0.9]$ with step size of $0.1$, $c\in\{0.1,0.5,0.9,0.999,0.99999\}$, convergence threshold for fixed point iteration $\in\{10^{-4},10^{-5},10^{-6}\}$, $d\in\{128,256,512,1024,2048\}$, and $M\in\{8,16,32,64\}$.
@@ 4.1.3 Results
Table 2 shows node classification results on the heterophilic datasets. Results marked with * in Table 2 are obtained from Platonov et al. (2023b) and in Table 3 from He et al. (2021). Applying ReP shows performance improvement for all backbones across most heterophilic datasets, with the most significant and consistent improvement observed in GCN. GCN+ReP achieves state-of-the-art performance in four out of seven datasets and the second-best performance in one. In the datasets where GCN+ReP attained state-of-the-art performance, the number of reverse layers was consistently above 64. This observation shows that deep layers of GNNs mjwith ReP can achieve superior performance without over-smoothing.
Table 3 shows node classification performance on the homophilic datasets. Due to spacing, we only report the results on three datasets. All results are reported in Appendix D. No significant performance changes were observed when ReP applied on homophilic node classification. The results confirm that the sharpened representations do not harm the prediction performance for homophily datasets where the forward aggregation is sufficient.
@@
The prediction performances with varying numbers of layers are reported in Figure 2. In both datasets, the prediction performance keeps increasing as the number of reverse layers increases. These results indicate that the reverse process is capable of deep stacking without suffering from over-smoothing, enabling the models to capture long-range dependencies effectively, which is known to be important, especially in heterophilic graphs. The prediction performance also tends to increase as we increase the number of forward steps up to 1024 in the Chameleon dataset.
Figure 2: Prediction performance with varying numbers of forward and reverse layers. We vary the number of layers in one direction. Due to memory constraints, we restricted the reverse depth used in Minesweeper to 256 or less.
@@
We evaluate whether the proposed reverse diffusion mitigates the over-smoothing issue. To measure the degree of over-smoothing, we adopt Graph Smoothness Level (GSL) proposed by Zhang et al. (2022) defined as:
$\operatorname{GSL}(\mathbf{X})=\frac{1}{N(N-1)}\sum_{i\in\mathcal{V}}\sum_{j\in\mathcal{V},j\neq i}\frac{\mathbf{x}_{i}\cdot\mathbf{x}_{j}}{\lVert{\mathbf{x}_{i}}\rVert_{F}\lVert{\mathbf{x}_{j}}\rVert_{F}},\;$ (7)
@@
Figure 3: Over smoothing levels measured by GSL over the number of layers. Rep (forward) denotes the measured GSL in the forward process of GCN+ReP. We compare the results with GCN of three different numbers of layers: 16, 32, 64.
@@
To validate that the reverse process produces a sharpened representation, we visualize label predictions on the Minesweeper dataset. The Minesweeper dataset is a binary classification task on a grid-structured graph, where the node with a positive label indicates the location of a mine. Each node, unless located on the boundaries, is connected to eight adjacent nodes, including the ones in the diagonal directions. The node feature is initialized with the number of mines in the adjacent nodes, and a one-hot representation of the feature is used as an initial representation of the node for learning.
@@
Based on the representations obtained from the forward and reverse processes, we trained a GCN+ReP model with a single-layer MLP as a prediction head. In Figure 4, we visualize the prediction results of two randomly sampled $7\times 7$ sub-grids from $100\times 100$ grid structure. For each example, we visualize the prediction results with node representations from 1) forward process, 2) reverse process, 3) both directions and 4) true labels displayed from upper left to lower right. Since the prediction head needs concatenated representations for prediction, to visualize the prediction results focused on a forward or reverse representation, we set the other node representation to be zero, e.g., to predict the mine using node representation at layer $\ell>0$, $(\mathbf{X}^{(\ell)}\|\mathbf{0})$ is fed into the prediction head.
[…]
@@
Figure 4: Visualization of node prediction on the minesweeper dataset. We visualize the prediction from 1) forward, 2) reverse, and 3) both representations, along with the ground truth labels.
@@
In both examples, the prediction results from the reverse process appear sharper while those from the forward process tend to be smooth. Additionally, the sharp prediction created by the reverse process significantly contributes to label prediction. Although in some cases, the reverse process can perfectly classify the location of mines, e.g., example 1, the other cases require representations from both directions to classify the mines correctly, e.g., example 2.
Figure 1 shows the changes in predictions over the number of layers. The darker the cell, the higher the predicted probability of a mine being present. As expected, the predictions and representations therein tend to be sharp as the number of reverse layers increases.
4.2 Prediction with Low Label Rate
We compare node classification performance in citation graphs with low label rates, used to mimic the long-range dependencies between nodes. According to Zhao & Akoglu (2020), typical datasets are generally sufficient for shallow architectures, making it difficult to identify the advantages of deep GNNs. As explained by Sun et al. (2020), in scenarios with a low training label rate, more layers are necessary to access distant supervision.
We perform node label classification tasks on the Cora, Citeseer, and Pubmed datasets, with 20 labels per class in the training set. The validation and test sets were consistently set to 500 and 1,000, respectively. Each experiment was conducted across ten random splits of the dataset, with the mean and standard deviation of the results being reported.
We compare node classification accuracy on three citation datasets, Cora, CiteSeer, and PubMed, with baseline methods including GCN and GRAND++ (Thorpe et al., 2022) which is proposed especially for low-labeling rate problems.
Table 4 shows the average accuracy and standard deviation. We find that performance improves when ReP applied. The reverse process is even more effective than GRAND++, a model tailored for the low-rate label prediction task.
Cora CiteSeer PubMed
GRAND++ 72.18±3.10 64.31±1.91 75.94±2.50
GCN 76.50±1.77 61.67±3.30 76.25±2.36
GCN+ReP 79.06±1.54 66.46±1.82 76.92±1.34
Table 4: Prediction accuracy on the citation graphs with 20-label-per-class.
@@ 5 Conclusion
To ensure invertibility, the Lipschitz constant of the forward process must be restricted to less than one, and the hidden dimension of weight parameters must remain constant. These restrictions limit the representation power and design choices within our framework. Investigating less restrictive invertible forms could lead to performance improvements.
@@ Appendix A Complexity Analysis
The method for calculating the forward process and the reverse process, when applied to a GCN with residual connections and weight sharing, is as follows:
$\mathbf{X}^{(\ell+1)}=\mathbf{X}^{(\ell)}+\hat{\mathbf{A}}\mathbf{X}^{(\ell)}\mathbf{W},$ (forward process) (8)
$\mathbf{X}^{(\ell-1)}=\mathbf{X}^{(\ell)}+\sum_{k=1}^{K}(-1)^{k}\hat{\mathbf{A}}^{k}\mathbf{X}^{(\ell)}\mathbf{W}^{k},$ (reverse process) (9)
For simplicity, we ignore an activation function.
@@ A.1 Forward Pass of Forward Process
Equation 8 involves three key operations: matrix multiplication between $\mathbf{X}$ and $\mathbf{W}$ with complexity $O(|\mathcal{V}|d^{2})$, matrix multiplication between sparse matrix $\hat{\mathbf{A}}$ and $\mathbf{X}^{(\ell)}\mathbf{W}$ with complexity $O(|\mathcal{E}|d)$, and matrix addition with complexity $O(|\mathcal{V}|d)$. Overall, the time complexity for one-layer is $O(|\mathcal{V}|d^{2}+|\mathcal{E}|d)$. This results in a total time complexity of $O(L_{D}|\mathcal{V}|d^{2}+L_{D}|\mathcal{E}|d)$ over $L_{D}$ layers. Memory cost is calculated as $O(L_{D}|\mathcal{V}|d+|\mathcal{E}|+d^{2})$.
A.2 Backward Pass of Reverse Process
@@
The time complexity of the backward pass primarily hinges on the computation of $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$. By applying the chain rule, this can be expressed as:
@@
$\frac{\partial\mathcal{L}}{\partial\mathbf{W}}=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}\frac{\partial\mathbf{X}^{(L_{D})}}{\partial\mathbf{X}^{(L_{D}-1)}}\cdots\frac{\partial\mathbf{X}^{(2)}}{\partial\mathbf{X}^{(1)}}\frac{\partial\mathbf{X}^{(1)}}{\partial\mathbf{W}}.$ (10)
Given $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}$, $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell-1)}}$ can be derived as follows:
$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell-1)}}=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}\frac{\partial\mathbf{X}^{(\ell)}}{\partial\mathbf{X}^{(\ell-1)}}$ $=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}\left(\frac{\partial}{\partial\mathbf{X}^{(\ell-1)}}\mathbf{I}\mathbf{X}^{(\ell-1)}+\frac{\partial}{\partial\mathbf{X}^{(\ell-1)}}\hat{\mathbf{A}}\mathbf{X}^{(\ell-1)}\mathbf{W}\right)$
$=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}+\hat{\mathbf{A}}^{\top}\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}\mathbf{W}^{\top}.$ (11)
And we can compute $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D}-1)}}$ as follows:
$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}\frac{\partial\mathbf{X}^{(L_{D})}}{\partial\mathbf{X}^{(L_{D}-1)}}$ $=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}\left(\frac{\partial}{\partial\mathbf{X}^{(L_{D}-1)}}\mathbf{I}\mathbf{X}^{(L_{D}-1)}+\frac{\partial}{\partial\mathbf{X}^{(L_{D}-1)}}\hat{\mathbf{A}}\mathbf{X}^{(L_{D}-1)}\mathbf{W}\right)$
$=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}+\hat{\mathbf{A}}^{\top}\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}\mathbf{W}^{\top}.$ (12)
By recursively applying Equation 11 to Equation 12, we can obtain $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(1)}}$ as follows:
@@
$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}\frac{\partial\mathbf{X}^{(L_{D})}}{\partial\mathbf{X}^{(L_{D}-1)}}\cdots\frac{\partial\mathbf{X}^{(2)}}{\partial\mathbf{X}^{(1)}}=\sum_{\ell=0}^{L_{D}-1}\binom{L_{D}-1}{\ell}({\hat{\mathbf{A}}^{\top}})^{\ell}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}\right)({\mathbf{W}^{\top}})^{\ell}.$ (13)
@@
Finally, $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ can be derived as the multiplication of $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(1)}}$ and $\frac{\partial\mathbf{X}^{(1)}}{\partial\mathbf{W}}$:
@@
$\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ $=\bigg{(}\sum_{\ell=0}^{L_{D}-1}\binom{L_{D}-1}{\ell}({\hat{\mathbf{A}}^{\top}})^{\ell}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}\right)({\mathbf{W}^{\top}})^{\ell}\bigg{)}\left(\frac{\partial}{\partial\mathbf{W}}\hat{\mathbf{A}}\mathbf{X}^{(0)}\mathbf{W}\right)$
@@
$=\left(\hat{\mathbf{A}}\mathbf{X}^{(0)}\right)^{\top}\bigg{(}\sum_{\ell=0}^{L_{D}-1}\binom{L_{D}-1}{\ell}({\hat{\mathbf{A}}^{\top}})^{\ell}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{D})}}\right)({\mathbf{W}^{\top}})^{\ell}\bigg{)}.$ (14)
@@
Note that $\hat{\mathbf{A}}$ is the sparse matrix, so matrix multiplication can be performed in $O(|\mathcal{E}|d)$. So, the time complexity of the term in the summation is $O(|\mathcal{V}|d^{2}+\ell d^{3}+\ell|\mathcal{E}|d)$. Thus, the time complexity for computing $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ is $O(L_{D}|\mathcal{V}|d^{2}+L_{D}^{2}d^{3}+L_{D}^{2}|\mathcal{E}|d)$, with a memory cost of $O(|\mathcal{V}|d+|\mathcal{E}|+d^{2})$.

### v2

@@ 4 Experiments
The experimental section focuses on validating two research questions: 1) Can the reverse process produce distinguishable representations? 2) Does the reverse process alleviate over-smoothing problems, enabling the construction of deeper layers?
Throughout this section, we denote models with additional reverse layers by ReP (Reverse Process). For example, GCN$+$ReP indicates the GCN backbone with the reverse process. We adopt weight sharing approach of GRAND-l for all experiments using ReP.
@@ 4.1.2 Experimental Setup and Baselines
We benchmark several neural architectures as baselines, including classic GNN models like GCN (Kipf & Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), and Graph Transformer (GT) (Shi et al., 2020) for more complex attention mechanisms. These baselines are augmented with skip connections and layer normalization. In addition, modifications proposed in Zhu et al. (2020) are made to GAT and GT, resulting in GAT-sep and GT-sep models. For heterophily-specific models, we use 10 models including H2GCN (Zhu et al., 2020), CPGNN (Zhu et al., 2021), GPR-GNN (Chien et al., 2020), FSGNN (Maurya et al., 2022), GloGNN (Li et al., 2022), FAGCN (Bo et al., 2021), GBK-GNN (Du et al., 2022), JacobiConv (Wang & Zhang, 2022), LRGNN (Liang et al., 2023), Ordered GNN (Song et al., 2023), ACM-GCN (Luan et al., 2022), and Dir-GNN (Rossi et al., 2023).
@@
For all experiments, we set the number of epochs to 1,000 and apply early stopping when there is no performance improvement for 100 consecutive epochs. For GRAND+ReP, we validate the hyperparameters that maximize the validation metric in the following ranges: learning rate $\in[10^{-5},10^{-1}]$, $T_{F},T_{R}\in[0,10]$, $d\in\{16,32,64,128,256,512\}$, $K\in[1,8]$, $d^{\prime}\in\{4,8,16,32,64,128\}$. With the Euler or Runge-Kutta methods, we search the step size over $[0.5,8]$ and tolerance scale over $[1,20000]$ with the Dormand-Prince method. For GCN+ReP and GAT+ReP, we validate the hyperparameters in the following ranges: learning rate$\in[10^{-5},10^{-1}]$, the number of forward and reverse layers $L_{F},L_{R}\in\{1,2,4,8,16,32,64,128,256,512\}$, dropout probability $\in[0,0.9]$ with step size of $0.1$, $c\in\{0.1,0.5,0.9,0.999,0.99999\}$, convergence threshold for fixed point iteration $\in\{10^{-4},10^{-5},10^{-6}\}$, $d\in\{128,256,512,1024,2048\}$, and $M\in\{8,16,32,64\}$. We fix the non-linear activation function to ReLU.
@@ 4.1.3 Results
Table 2 shows node classification results on the heterophilic datasets. Results marked with * and in Table 2 are obtained from Platonov et al. (2023b) and Rossi et al. (2023), and * in Table 3 from He et al. (2021). Applying ReP shows performance improvement for all backbones across most heterophilic datasets, with the most significant and consistent improvement observed in GCN. GCN+ReP achieves state-of-the-art performance in four out of seven datasets and the second-best performance in one. In the datasets where GCN+ReP attained state-of-the-art performance, the number of reverse layers was consistently above 64. This observation shows that deep layers of GNNs with ReP can achieve superior performance without over-smoothing.
Table 3 shows node classification performance on the homophilic datasets. Due to spacing, we only report the results on three datasets. All results are reported in Appendix D. No significant performance changes were observed when ReP applied on homophilic node classification. The results confirm that the distinguishable representations do not harm the prediction performance for homophily datasets where the forward aggregation is sufficient.
@@
The prediction performances with varying numbers of layers are reported in Figure 2. In both datasets, the prediction performance keeps increasing as the number of reverse layers increases. These results indicate that the reverse process is capable of deep stacking to mitigate over-smoothing. This enables the models to capture long-range dependencies effectively, which is known to be important, especially in heterophilic graphs. The prediction performance also tends to increase as we increase the number of forward steps up to 1024 in the Chameleon dataset.
Figure 2: Prediction performance with varying the number of forward and reverse layers. We vary the number of layers (depth) in one direction. Due to memory constraints, we restricted the reverse depth used in Minesweeper to 256 or less.
@@
We evaluate whether the proposed reverse process mitigates the over-smoothing issue. To measure the degree of over-smoothing, we adopt Graph Smoothness Level (GSL) proposed by Zhang et al. (2022) defined as:
$\operatorname{GSL}(\mathbf{X})=\frac{1}{|\mathcal{V}|(|\mathcal{V}|-1)}\sum_{i\in\mathcal{V}}\sum_{j\in\mathcal{V},j\neq i}\frac{\mathbf{x}_{i}\cdot\mathbf{x}_{j}}{\lVert{\mathbf{x}_{i}}\rVert_{F}\lVert{\mathbf{x}_{j}}\rVert_{F}},\;$ (9)
@@
Figure 3: Over-smoothing levels measured by GSL over the number of layers (depth). ReP (forward) denotes the measured GSL in the forward process of GCN+ReP. We compare the results with GCN of three different depths: 16, 32, 64.
@@
To validate that the reverse process produces a distinguishable representation, we visualize label predictions on the Minesweeper dataset. The Minesweeper dataset is a binary classification task on a grid-structured graph, where the node with a positive label indicates the location of a mine. Each node, unless located on the boundaries, is connected to eight adjacent nodes, including the ones in the diagonal directions. The node feature is initialized with the number of mines in the adjacent nodes, and a one-hot representation of the feature is used as an initial representation of the node for learning.
@@
Based on the representations obtained from the forward and reverse processes, we trained a GCN+ReP model with a single-layer MLP as a prediction head. In Figure 4, we visualize the prediction results of two randomly sampled $7\times 7$ sub-grids from $100\times 100$ grid structure. For each example, we visualize the prediction results with node representations from 1) forward process, 2) reverse process, and 3) both directions, as well as 4) the true labels, displayed from upper left to lower right. In the visualization of the true labels, black cells indicate the presence of mine, and white cells indicate its absence. In the visualization of the prediction results, the darker the cell, the higher the predicted probability of a mine being present. Since the prediction head needs concatenated representations for prediction, to visualize the prediction results focused on a forward or reverse representation, we set the other node representation to be zero, e.g., to predict the mine using node representation at layer $\ell>0$, $(\mathbf{X}^{(\ell)}\|\mathbf{0})$ is fed into the prediction head.
(a) Example 1
(b) Example 2
@@
Figure 4: Visualization of node prediction on the Minesweeper dataset. We visualize the prediction from 1) forward, 2) reverse, and 3) both representations, along with 4) the ground truth labels.
@@
In both examples, the prediction results from the reverse process appear distinguishable, while those from the forward process tend to be smooth. Additionally, the distinguishable prediction created by the reverse process significantly contributes to label prediction. Although in some cases, the reverse process can perfectly classify the location of mines, e.g., Example 1, the other cases require representations from both directions to classify the mines correctly, e.g., Example 2.
Figure 1 shows the changes in predictions over the number of layers on the $10\times 10$ sampled sub-grids. We follow the same visualization procedure with Figure 4. As expected, the predictions and representations tend to be more distinguishable as the number of reverse layers increases.
4.2 Inversion Error
Although the invertibility of Algorithm 1 is guaranteed in theory, the inversion may not be achieved due to numerical errors. To verify the fixed point method, we conduct the experiment to restore the inputs from the outputs of GCN and GAT at a depth of $64$. We set the scaling coefficient $c$ to $0.99999$ and the number of fixed point iterations to eight, which is challenging due to the large coefficient (note that the Lipschitz of 1 is non-invertible) and the small number of iterations. Table 4 shows the mean absolute error between the original inputs and restored input data. As the results show, the inversion error is negligible in practice.
GCN+ReP GAT+ReP
Squirrel $3.36\times 10^{-5}$ $3.28\times 10^{-5}$
Chameleon $2.23\times 10^{-5}$ $2.41\times 10^{-5}$
Roman-empire $2.79\times 10^{-5}$ $3.56\times 10^{-5}$
Amazon-ratings $4.31\times 10^{-5}$ $3.46\times 10^{-5}$
Table 4: Mean absolute error between the original inputs and restored input data by Algorithm 1.
@@ 5 Conclusion
To ensure invertibility, the Lipschitz constant of the forward process must be restricted, and the hidden dimension of weight parameters must remain constant. These restrictions limit the representation power and design choices. Investigating less restrictive invertible forms could lead to performance improvements.
@@
Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (RS-2019-II191906, Artificial Intelligence Graduate School Program(POSTECH)) and supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT)(RS-2024-00337955 and RS-2023-00217286).
@@ Appendix A Complexity Analysis
The forward and reverse processes of a GCN with residual connections and weight sharing are as follows:
[…]
where $M$ is the number of fixed point iterations. For simplicity, we ignore an activation function.
@@ A.1 Forward Pass of Forward Process
Equation 10 involves three key operations: matrix multiplication between $\mathbf{X}$ and $\mathbf{W}$ with complexity $O(|\mathcal{V}|d^{2})$, matrix multiplication between sparse matrix $\hat{\mathbf{A}}$ and $\mathbf{X}^{(\ell)}\mathbf{W}$ with complexity $O(|\mathcal{E}|d)$, and matrix addition with complexity $O(|\mathcal{V}|d)$. Overall, the time complexity for a single layer is $O(|\mathcal{V}|d^{2}+|\mathcal{E}|d)$. Therefore, the total time complexity over $L_{F}$ forward layers is $O(L_{F}|\mathcal{V}|d^{2}+L_{F}|\mathcal{E}|d)$. Memory cost is calculated as $O(L_{F}|\mathcal{V}|d+|\mathcal{E}|+d^{2})$.
A.2 Backward Pass of Forward Process
@@
The time complexity of the backward pass is primarily determined by the computation cost of $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$. By applying the chain rule, the gradient can be expressed as:
$\frac{\partial\mathcal{L}}{\partial\mathbf{W}}=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}\frac{\partial\mathbf{X}^{(L_{F})}}{\partial\mathbf{X}^{(L_{F}-1)}}\cdots\frac{\partial\mathbf{X}^{(2)}}{\partial\mathbf{X}^{(1)}}\frac{\partial\mathbf{X}^{(1)}}{\partial\mathbf{W}}.$ (12)
@@
By sequentially multiplying $\frac{\partial\mathbf{X}^{(n)}}{\partial\mathbf{X}^{(n-1)}}$ for $n=L_{F},\cdots,2$ on the right side of $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}$, we derive:
@@
$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}\frac{\partial\mathbf{X}^{(L_{F})}}{\partial\mathbf{X}^{(L_{F}-1)}}\cdots\frac{\partial\mathbf{X}^{(2)}}{\partial\mathbf{X}^{(1)}}=\sum_{\ell=0}^{L_{F}-1}\binom{L_{F}-1}{\ell}({\hat{\mathbf{A}}^{\top}})^{\ell}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}\right)({\mathbf{W}^{\top}})^{\ell}\;.$ (13)
To demonstrate Equation 13, we illustrate part of the sequential multiplication process. Multiplying $\frac{\partial\mathbf{X}^{(L_{F})}}{\partial\mathbf{X}^{(L_{F}-1)}}$ to the right side of $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}$, we get:
$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}\frac{\partial\mathbf{X}^{(L_{F})}}{\partial\mathbf{X}^{(L_{F}-1)}}$ $=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}\left(\frac{\partial}{\partial\mathbf{X}^{(L_{F}-1)}}\mathbf{I}\mathbf{X}^{(L_{F}-1)}+\frac{\partial}{\partial\mathbf{X}^{(L_{F}-1)}}\hat{\mathbf{A}}\mathbf{X}^{(L_{F}-1)}\mathbf{W}\right)$
$=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}+\hat{\mathbf{A}}^{\top}\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}\mathbf{W}^{\top}.$ (14)
Next, we multiply $\frac{\partial\mathbf{X}^{(L_{F}-1)}}{\partial\mathbf{X}^{(L_{F}-2)}}$ on the right side of Equation 14, yielding:
$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}\frac{\partial\mathbf{X}^{(L_{F})}}{\partial\mathbf{X}^{(L_{F}-1)}}\frac{\partial\mathbf{X}^{(L_{F}-1)}}{\partial\mathbf{X}^{(L_{F}-2)}}$ $=\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}+\hat{\mathbf{A}}^{\top}\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}\mathbf{W}^{\top}\right)\left(\frac{\partial}{\partial\mathbf{X}^{(L_{F}-2)}}\mathbf{I}\mathbf{X}^{(L_{F}-2)}+\frac{\partial}{\partial\mathbf{X}^{(L_{F}-2)}}\hat{\mathbf{A}}\mathbf{X}^{(L_{F}-2)}\mathbf{W}\right)$
$=\sum_{\ell=0}^{2}\binom{2}{\ell}({\hat{\mathbf{A}}^{\top}})^{\ell}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}\right)({\mathbf{W}^{\top}})^{\ell}\;.$ (15)
Repeating the process above, we can derive Equation 13.
@@
Finally, $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ is derived by multiplying Equation 13 with $\frac{\partial\mathbf{X}^{(1)}}{\partial\mathbf{W}}$:
$\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ $=\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(1)}}\right)\left(\frac{\partial\mathbf{X}^{(1)}}{\partial\mathbf{W}}\right)$
@@
$=\bigg{(}\sum_{\ell=0}^{L_{F}-1}\binom{L_{F}-1}{\ell}({\hat{\mathbf{A}}^{\top}})^{\ell}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}\right)({\mathbf{W}^{\top}})^{\ell}\bigg{)}\left(\frac{\partial}{\partial\mathbf{W}}\hat{\mathbf{A}}\mathbf{X}^{(0)}\mathbf{W}\right)$
@@
$=\left(\hat{\mathbf{A}}\mathbf{X}^{(0)}\right)^{\top}\bigg{(}\sum_{\ell=0}^{L_{F}-1}\binom{L_{F}-1}{\ell}({\hat{\mathbf{A}}^{\top}})^{\ell}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(L_{F})}}\right)({\mathbf{W}^{\top}})^{\ell}\bigg{)}.$ (16)
@@
The time complexity of the term in the summation is $O(L_{F}+\ell|\mathcal{E}|d+|\mathcal{V}|d^{2}+\ell d^{3})$. Therefore, the overall time complexity for computing $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ is $O(L_{F}^{2}d^{3}+L_{F}|\mathcal{V}|d^{2}+L_{F}^{2}|\mathcal{E}|d)$, with a memory cost of $O(|\mathcal{V}|d+|\mathcal{E}|+d^{2})$.